OpenStudy (anonymous):

Solve the solution set using interval notation. 1/2 (2x+10)> 3/4 (x-1)

OpenStudy (anonymous):

\[x+5 > \frac{3}{4}x-\frac{3}{4}\] \[4x+20>3x-3\] \[x>-23\] \[(-23,\infty)\]